# Reliable algorithm selection for machine learning-guided design

### v1

@@ 4 Related work
Design algorithm selection belongs to a body of work on managing predictive uncertainty in machine learning-guided design. Design algorithms can be constrained to stay close to the training data or trusted reference points, such that predictions remain trustworthy [8, 27, 7, 21, 45], and out-of-distribution detection methods can be used to flag individual designs whose predictions are unreliable [12]. A variety of methods have been used to quantify predictive uncertainty for individual designs in biomolecular design tasks [19], including those based on ensembling [35, 20], Gaussian processes [23, 48], and evidential learning [41]. In particular, conformal prediction techniques construct prediction sets for designs that also have frequentist-style guarantees: the sets contain the design labels with guaranteed probability, where the probability is over drawing designs from the design distribution [15, 30]. However, it is unclear how to invoke such statements to make decisions about which designs to use, as there is no guarantee regarding the prediction sets for any specific designs of interest—for example, those whose prediction sets satisfy a desired condition.
Moreover, quantifying uncertainty for individual designs is perhaps unnecessary for the goal of many design campaigns in practice. Success often does not necessitate that every design performs as desired, but only that sufficiently many do so, regardless of which specific ones [52]. For example, Wang et al. [49] develop a method that selects designs from a pool of individual candidates, such that the selected subset contains a desired expected number whose label surpasses a threshold. However, the method requires held-out labeled data from the same distribution as the candidates, which is not always available in practice.
@@
Conformal selection methods [24, 25, 3] combine ideas from conformal prediction and multiple testing to also select designs from a candidate pool, with guaranteed upper bounds on the false discovery rate, or expected proportion of selected designs whose label does not surpass a desired threshold. The setting of these methods differs from ours in that they assume access to a pool of candidates that are exchangeable—for example, candidates drawn i.i.d. from some distribution. However, depending on the application, it may not always be straightforward to narrow down a large design space $\mathcal{X}$ to a suitable pool of candidates to begin with, without selecting a design algorithm configuration with which to generate designs. The goal of our work is therefore to choose from a set of candidate design algorithm configurations, each of which will produce a different distribution of designs when deployed. Our approach aims to select configurations that will achieve a population-level notion of success, which we formalize as a user-specified functional of the distribution of design labels, $\theta_{\lambda}$, surpassing some threshold. Examples of this functional include the mean design label and the exceedance, or the fraction of design labels that exceed a threshold, and more generally any population-level quantity that can be expressed as minimizer of a convex loss (Appendix B.1).
@@
Bayesian optimization (BO) [37] is a well-studied paradigm for iteratively selecting designs, acquiring their labels, and updating a predictive model in order to optimize a property of interest. In each round, in the language of our framework, BO prescribes a design algorithm: it chooses the design (or, more similar to our setting, the batch of designs [13]) that globally maximizes some acquisition function that quantifies desirability based on the model’s predictions, and which typically incorporates the model’s predictive uncertainty, such as the expected improvement of a design’s label over the best one found thus far. A typical goal of BO algorithms is to converge to the global optimum as the rounds progress indefinitely, under various regularity conditions on the property one seeks to optimize [43, 11, 5]. In contrast, the goal of design algorithm selection is to achieve guarantees on the distribution of design labels to be imminently proposed. Such guarantees are particularly useful and can help justify designs to stakeholders when acquiring labels for even one round is time- or resource-intensive, and the priority is to find designs that achieve specific improvements within one or a few rounds, rather than to eventually find the best possible design as the number of rounds goes to infinity. However, nothing precludes batch BO algorithms from being used as design algorithms within our framework if one deems them promising for the design task at hand. Configurations of batch BO algorithms with different hyperparameter settings or acquisition functions, for example, can be included on the menu alongside any other design algorithm configuration under consideration.
@@
Similar to our work, Wheelock et al. [52] also focus on characterizing the design label distribution. After constructing a forecast, or model of the conditional distribution of the label, for every design, their approach uses an equally weighted mixture of designs’ forecasts as a model of the design label distribution. A rich body of work also exists on calibrating forecasts, such that various aspects of these conditional distributions are statistically consistent with held-out labeled data [17, 26, 42]. Our approach differs in that it directly estimates the functional of the design label distribution, $\theta_{\lambda}$, that determines whether a configuration is successful, rather than characterizing predictive uncertainty for individual designs. That is, our method handles prediction error in a way that is specifically tailored for the endpoint of selecting which design algorithm configuration to deploy.
@@
At a technical level, our work uses prediction-powered inference techniques [2] adapted for covariate shift [38], in order to conduct statistically valid hypothesis tests of whether configurations are successful. The multiple hypothesis testing approach is similar to that of Angelopoulos et al. [1], who use multiple testing to select hyperparameter settings for predictive models in order to achieve desired risk values with high-probability guarantees.
@@ 5 Experiments
We first demonstrate that our method selects successful configurations with high probability, as guaranteed by theory, when the design and labeled densities are known. Next, we show that it still selects successful configurations more effectively than alternative methods when these densities are estimated.
Two metrics are of interest: error rate and selection rate. Error rate is the empirical frequency at which a method incorrectly selects a configuration that fails the success criterion (Eq. 1), computed over multiple trials of sampling designs from each configuration as well as held-out labeled data (for methods that require it). Selection rate is the empirical frequency over those same trials at which a method selects anything at all. A good method achieves a low error rate while maintaining a high selection rate, which may be challenging for ambitious success criteria.
@@
Gaussian-mixture-model forecast method
For any sequence, Wheelock et al. [52] model the conditional distribution of the label as a mixture of two Gaussians with sequence-specific parameters, which capture beliefs over the label if the sequence is “functional” or “nonfunctional.” After training this model, they infer forecasts for every design produced by a configuration $\lambda$, where a hyperparameter $q\in[0,1]$ controls, roughly speaking, how much the forecasts deviate from the training data (see Appendix C.2 for details). The equally weighted mixture of these forecasts serves as a model, $P^{\text{GMM}}_{\lambda}$, of the design label distribution for configuration $\lambda$. We select configurations using the following procedure, called GMMForecasts : for each $\lambda\in\Lambda$, we select $\lambda$ if $\mathbb{E}_{Y\sim P^{\text{GMM}}_{\lambda}}[g(Y)]\geq\tau$.
Calibrated forecast method
This method similarly models the design label distribution as the equally weighted mixture of forecasts for the designs produced by a configuration. Each forecast is initially modeled as a Gaussian with mean and variance set to the predictive mean and variance, respectively, given by the predictive model. We then use the labeled data to calibrate these forecasts post hoc using isotonic regression [26] (see Appendix C.3 for details). We select configuration $\lambda$ if $\mathbb{E}_{Y\sim P^{\text{cal}}_{\lambda}}[g(Y)]\geq\tau$, where $P^{\text{cal}}_{\lambda}$ is the equally weighted mixture of the calibrated forecasts for the designs. This procedure is called CalibratedForecasts .
@@
We adapted conformal prediction techniques to conduct design algorithm selection, as detailed in Appendix C.4. This method has similar theoretical guarantees as our method, but is prohibitively conservative: across all our experiments, it achieved an error rate of zero by never selecting anything. We therefore exclude these results for clarity of visualization in Figures 3, 4, and 6.
Figure 2: Library design for protein GB1. Mean prediction and KL divergence from the training distribution of the $101$ design algorithm configurations on the menu. Each dot corresponds to the design distribution, $q_{\lambda}$, for a specific value of the temperature hyperparameter, $\lambda$. Two red-outlined squares correspond to design distributions for a low temperature ($\lambda=0.2$) and a high temperature ($\lambda=0.65$), whose parameters are shown in the insets (purple and green heatmaps, respectively).
@@ 5.1 Algorithm selection for designing protein GB1
The design task for the first set of experiments was to design novel protein sequences that have high binding affinity to an immunoglobulin, by specifying the amino acids at four particular sites of a protein called GB1. These experiments simulate library design, an important practical setting in which both the design and labeled densities have closed-form expressions. Specifically, the most time- and cost-effective protocols today for synthesizing protein sequences in the wet lab can be described mathematically as sampling the amino acid at each site independently from a site-specific categorical distribution, whose parameters we can set; the density for any sequence is then the product of the probabilities of the amino acid at each site. We follow the design algorithm developed by Zhu, Brookes, & Busia et al. (2024): we first train a predictive model of binding affinity (Appendix D), $f$, and then set the parameters of the site-specific categorical distributions such that sequences with high predictions have high likelihood, as follows.
@@
Figure 3: Design algorithm selection for designing protein GB1. (a) Error rate (top; lower is better) and selection rate (bottom; higher is better) for all methods, for range of values of $\tau$, the desired mean design label. GMMForecasts with hyperparameter $q\in\{0,0.5\}$ never selected anything, resulting in error and selection rates of zero for all $\tau$. (b) For each method, the median (solid line) and $20^{\text{th}}$ to $80^{\text{th}}$ percentiles (shaded regions) of the lowest mean design label of selected configurations, over trials for the method did not return the empty set. Dots mark where each median trajectory ends (i.e., the value of $\tau$ beyond which a method ceases to select any configuration, and the mean design label achieved for that $\tau$).
@@
To simulate training and held-out labeled data, we sampled sequences from a common baseline called the NNK library, which is close to uniform categorical distributions at every site, but slightly modified to reduce the probability of stop codons. For the labels, we used a data set that contains experimentally measured binding affinities for every sequence in $\mathcal{X}$ [53]—that is, all $20^{4}$ variants of protein GB1 at 4 specific sites. These labels were log-ratios relative to wild type GB1, such that the wild type’s label is 00.
@@
The menu contained $101$ values of $\lambda$ between $0.2$ and $0.7$. We used the following success criteria: that the mean design label surpasses $\tau$, for $\tau\in[0,1.5]$ (for comparison, the mean training label was $-4.8$), and that the exceedance over $1$ (i.e., the fraction of design labels that exceed $1$) surpasses $\tau$, for $\tau\in[0,1]$ (for comparison, the training labels’ exceedance over $1$ was $0.006$).
@@
For any $\lambda\in\Lambda$, the true value of $\theta_{\lambda}$ has a closed-form expression, since we have labels for all sequences in $\mathcal{X}$ [53]. We then computed the error rate for each method as $\sum_{t=1}^{T}\mathbbm{1}[\exists\lambda\in\hat{\Lambda}\text{ s.t. }\theta_{\lambda}<\tau]/T$. We used $\alpha=0.1$ as a representative value for the desired error rate.
@@ Results
The prediction-only and CalibratedForecasts methods had error rates of 100% for most values of $\tau$ (Fig. 3a). Furthermore, particularly for the former, the true mean design label produced with selected temperatures could be considerably lower than $\tau$ (Fig. 3b).
Our method’s error rate was below $\alpha$ for all values of $\tau$ (Fig. 3a). That is, with frequency at least $1-\alpha$, our method selected temperatures resulting in mean design labels that surpassed $\tau$ (Fig. 3b). The selection rate was also 100% for many values of $\tau$, though it gradually declined for $\tau>1$.
For GMMForecasts , note that smaller values of the hyperparameter $q\in[0,1]$ yield forecasts that are, roughly speaking, more similar to the training data. Using $q\in\{0,0.5\}$ was prohibitively conservative and never selected anything on any trial. Using the maximum value, $q=1$, did result in high selection rates, but only for $\tau<0.5$ (Fig. 3a). That is, if a more ambitious mean label was requested, the method incorrectly reported that no temperature could achieve it.
Figure 4: Design algorithm selection for designing RNA binders. (a) Error rate (top; lower is better) and selection rate (bottom; higher is better) for all methods, for range of values of $\tau$, the desired mean design label. (b) For each method, the median (solid line) and $20^{\text{th}}$ to $80^{\text{th}}$ percentiles (shaded regions) of the lowest mean design label of selected configurations, over trials for the method did not return the empty set. Dots mark where each median trajectory ends (i.e., the value of $\tau$ beyond which a method ceases to select any configuration, and the mean design label achieved for that $\tau$).
@@ 5.2 Algorithm selection for designing RNA binders
These next experiments show the utility of our method in a setting that involves a variety of design algorithms and predictive models on the menu, and that requires density ratio estimation. The task is to design length-$50$ RNA sequences that bind well to an RNA target, where the label is the ViennaFold binding energy [28].
@@
To simulate training and held-out labeled data, we generated random mutants of a “seed” sequence, with $0.08$ probability of mutation at each site. Each mutant, $x$, was given a noisy label, $y=\textsc{BindingEnergy}(x)+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma=0.02)$.
@@
The menu contains the following design algorithms and respective hyperparameter settings (see Appendix E.1 and Fig. 7 for details): AdaLead [40], with its threshold hyperparameter, $\kappa$, set to values in $[0.2,0.01]$; the algorithm used by Biswas et al. [7], which approximately runs MCMC sampling from $p(x)\propto\exp(f(x)/T)$, where we set the temperature hyperparameter $T$ to values in $[0.005,0.02]$; Conditioning by Adaptive Sampling [8] and Design by Adaptive Sampling [9], with their quantile hyperparameter, $Q$, set to values in $[0.1,0.9]$; and Proximal Exploration [31] with default hyperparameters, as the authors did not note any critical hyperparameters. Each of these design algorithms and respective hyperparameter settings was run with three different predictive models, resulting in a menu of $78$ configurations: a ridge regression model, with the regularization hyperparameter set by leave-one-out cross-validation, an ensemble of feedforward neural networks, and an ensemble of convolutional neural networks. The selection criterion was that the mean design label surpasses $\tau$, for $\tau\in[0.28,0.5]$ (for comparison, the mean training label was $0.28$).
@@
None of the configurations have a closed-form design density. We therefore used multinomial logistic regression-based density ratio estimation (MDRE) [44], which trains a classifier between designs and labeled sequences to estimate the density ratios between their distributions (see Appendix E.2 for details).

### v2

@@ 4 Related Work
Design algorithm selection belongs to a body of work on managing predictive uncertainty in machine learning-guided design. Design algorithms can be constrained to stay close to the training data or trusted reference points, such that predictions remain trustworthy (Brookes et al.,, 2019; Linder et al.,, 2020; Biswas et al.,, 2021; Gruver et al.,, 2023; Tagasovska et al.,, 2024), and out-of-distribution detection methods can flag individual designs whose predictions are unreliable (Damani et al.,, 2023). A variety of methods have been used to quantify predictive uncertainty for individual designs in biomolecular design tasks (Greenman et al.,, 2025), including those based on ensembling (Scalia et al.,, 2020; Gruver et al.,, 2021), Gaussian processes (Hie et al.,, 2020; Tran et al.,, 2020), and evidential learning (Soleimany et al.,, 2021). In particular, conformal prediction techniques produce prediction sets for designs that also have frequentist-style guarantees: the sets contain the design labels with guaranteed probability, where the probability is over drawing designs from the design distribution (Fannjiang et al.,, 2022; Prinster et al.,, 2023). However, it is unclear how to invoke such statements to make decisions about which designs to use, as there is no guarantee regarding the prediction sets for any specific designs of interest—for example, those whose prediction sets satisfy a desired condition.
@@
Moreover, quantifying uncertainty for individual designs is perhaps unnecessary for the goal of many design campaigns in practice. Success often does not necessitate that every design performs as desired, but only that sufficiently many do so, regardless of which specific ones (Wheelock et al.,, 2022). For example, Wang et al., (2022) develop a method that selects designs from a pool of individual candidates, such that the selected subset contains a desired expected number whose label surpasses a threshold. Conformal selection methods (Jin and Candès,, 2022, 2023; Bai and Jin,, 2024) combine ideas from conformal prediction and multiple testing to also select designs from a pool of candidates, with guaranteed upper bounds on the false discovery rate, or expected proportion of selected designs whose label falls below a desired threshold. The setting of these methods differs from ours in that they assume access to a pool of candidates that are exchangeable—for example, candidates drawn i.i.d. from some distribution. However, depending on the application, it may not always be straightforward to narrow down a large design space $\mathcal{X}$ to a suitable pool of candidates to begin with, without selecting a design algorithm configuration with which to generate designs. The goal of our work is therefore to choose from a set of candidate design algorithm configurations, each of which will produce a different distribution of designs when deployed. Our approach aims to select configurations that will achieve a user-specified success criterion, which we formalize as a user-specified population-level metric, $\theta_{\lambda}$, surpassing some threshold.
@@
Bayesian optimization (BO) (Shahriari et al.,, 2016) is a well-studied paradigm for iteratively selecting designs, acquiring their labels, and updating a predictive model in order to optimize a property of interest. In each round, using the language of our framework, BO prescribes a design algorithm: it chooses the design (or batch of designs (Desautels et al.,, 2014)) that globally maximizes some acquisition function that quantifies desirability based on the model’s predictions, and which typically incorporates the model’s predictive uncertainty, such as the expected improvement of a design’s label over the best one found thus far. A typical goal of BO algorithms is to converge to the global optimum as the rounds progress, under regularity conditions on the property of interest (Srinivas et al.,, 2010; Bull,, 2011; Berkenkamp et al.,, 2019). In contrast, the goal of design algorithm selection is to achieve criteria on the distribution of design labels to be imminently proposed. Such guarantees can help justify designs to stakeholders when acquiring labels for even one round is resource-intensive, and the priority is to find designs that achieve specific improvements within one or a few rounds, rather than to eventually find the best possible design as the rounds progress indefinitely. However, nothing precludes batch BO algorithms from being used as design algorithms in our framework: configurations of these algorithms with different hyperparameter settings or acquisition functions, for example, can be on the menu.
@@
Similar to our work, Wheelock et al., (2022) also aim to characterize the design label distribution. To do so, their approach focuses on how to construct forecasts, or models of the conditional distribution of the label, for individual designs. An equally weighted mixture of designs’ forecasts then serves as a model of the design label distribution. A rich body of work also exists on calibrating forecasts, such that various aspects of these conditional distributions are statistically consistent with held-out labeled data (Gneiting et al.,, 2007; Kuleshov et al.,, 2018; Song et al.,, 2019). Our approach differs in that it directly estimates the population-level metric, $\theta_{\lambda}$, that determines whether a configuration is successful, rather than modeling the labels of individual designs. That is, our method handles prediction error in a way that specifically targets the endpoint of selecting which configuration to deploy.
@@
At a technical level, our work uses prediction-powered inference techniques (Angelopoulos et al.,, 2023) adapted for covariate shift, in order to conduct statistically valid hypothesis tests of whether configurations are successful. The multiple hypothesis testing approach is similar to that of Angelopoulos et al., (2021), who use multiple testing to set hyperparameters for predictive models to achieve desired risk values with high-probability guarantees.
@@ 5 Experiments
[…]
Two metrics are of interest: error rate and selection rate. Error rate is the empirical frequency at which a method selects a configuration that fails the success criterion (Eq. 1), over multiple trials of sampling held-out labeled data (for methods that require it) as well as designs from each configuration. Selection rate is the empirical frequency over those same trials at which a method selects anything at all. A good method achieves a low error rate while maintaining a high selection rate, which may be challenging for ambitious success criteria.
@@
Forecast-based methods ( GMMForecasts and CalibratedForecasts )
The next two methods both construct a forecast, or model of the conditional distribution of the label, for every design generated by a configuration $\lambda$. They then use the equally weighted mixture of the designs’ forecasts as a model of the design label distribution, $p_{Y;\lambda}$, and select $\lambda$ if this model satisfies the success criterion. The two methods differ in how they construct the forecast for a designed sequence, as follows.
GMMForecasts
follows Wheelock et al., (2022), who model the conditional distribution of the label as a mixture of two Gaussians with sequence-specific parameters, which capture beliefs over the label if the sequence is “functional” or “nonfunctional.” After training this model, forecasts are made for every design produced by a configuration $\lambda$, using different values of a hyperparameter $q\in[0,1]$ that controls, roughly speaking, how much the forecasts deviate from the training data (see Appendix C.2 for details). The equally weighted mixture of these forecasts, $P^{\text{GMM}}_{\lambda}$, serves as a model of the design label distribution for configuration $\lambda$. We select $\lambda$ if $\mathbb{E}_{Y\sim P^{\text{GMM}}_{\lambda}}[g(Y)]\geq\tau$.
For CalibratedForecasts , the forecast for each design is initially modeled as a Gaussian with mean and variance set to the predictive mean and variance, respectively, given by the predictive model. These forecasts are then transformed or calibrated using isotonic regression to be more consistent with the labeled data (Kuleshov et al.,, 2018) (see Appendix C.3 for details). We select configuration $\lambda$ if $\mathbb{E}_{Y\sim P^{\text{cal}}_{\lambda}}[g(Y)]\geq\tau$, where $P^{\text{cal}}_{\lambda}$ is the equally weighted mixture of the calibrated forecasts for designs from $\lambda$.
@@
We adapted conformal prediction techniques to conduct design algorithm selection (see Appendix C.4 for details). This method has similar theoretical guarantees to ours, but is prohibitively conservative: it never selected any configuration in our experiments. We therefore exclude these results for clarity of visualization.
@@ 5.1 Algorithm Selection for Designing Protein GB1
The design task for the first set of experiments was to design novel protein sequences that have high binding affinity to an immunoglobulin, by specifying the amino acids at four particular sites of a protein called GB1. These experiments simulate library design, an important practical setting in which both the design and labeled densities have closed-form expressions. Specifically, the most time- and cost-effective protocols today for synthesizing protein sequences in the wet lab can be described mathematically as sampling the amino acid at each site independently from a site-specific categorical distribution, whose parameters we can set; the density for any sequence is then the product of the probabilities of the amino acid at each site. We follow the design algorithm developed by Zhu, Brookes, & Busia et al. (2024): after training a predictive model of binding affinity (see Appendix D for details), $f$, we set the parameters of the site-specific categorical distributions such that sequences with high predictions have high likelihood, as follows.
@@
Figure 2: Library design for protein GB1. Mean prediction and KL divergence from the training distribution of the $101$ design algorithm configurations on the menu. Each dot corresponds to the design distribution, $q_{\lambda}$, for a specific value of the temperature hyperparameter, $\lambda$. Two red-outlined squares correspond to design distributions for the lowest temperature ($\lambda=0.2$) and the highest temperature ($\lambda=0.7$), whose parameters values are shown in the inset (top and bottom heatmaps, respectively).
Figure 3: Design algorithm selection for designing protein GB1. (a) Error rate (top; lower is better) and selection rate (bottom; higher is better) for all methods, for range of values of $\tau$, the desired mean design label. For reference, the mean label of the training data was $-4.8$. GMMForecasts with hyperparameter $q\in\{0,0.5\}$ (dark and medium purple lines) never selected anything, resulting in error and selection rates of zero for all $\tau$. (b) For each method, the median (solid line) and $20^{\text{th}}$ to $80^{\text{th}}$ percentiles (shaded regions) of the lowest mean design label achieved by selected configurations, over trials for which the method did not return the empty set. Dots mark where each median trajectory ends (i.e., the value of $\tau$ beyond which a method ceases to select any configuration, and the lowest mean design label of configurations selected for that $\tau$). Results on or above the dashed diagonal line indicate that selected configurations are successful.
Figure 4: Design algorithm selection for designing RNA binders. (a) Error rate (top; lower is better) and selection rate (bottom; higher is better) for all methods, for range of values of $\tau$, the desired mean design label. For reference, the mean training label was $0.28$. (b) For each method, the median (solid line) and $20^{\text{th}}$ to $80^{\text{th}}$ percentiles (shaded regions) of the lowest mean design label achieved by selected configurations, over trials for which the method did not return the empty set. Dots mark where each median trajectory ends (i.e., the value of $\tau$ beyond which a method ceases to select any configuration, and the lowest mean design label of configurations selected for that $\tau$). Results on or above the dashed diagonal line indicate that selected configurations are successful.
@@
To simulate training and held-out labeled data, we sampled sequences from the NNK library, a common baseline that is close to uniform categorical distributions at every site, but slightly modified to reduce the probability of stop codons. For the labels, we used a data set that contains experimentally measured binding affinities for all sequences $x\in\mathcal{X}$ (Wu et al.,, 2016)—that is, all $20^{4}$ variants of protein GB1 at 4 specific sites. Labels were log-ratios relative to the wild type, whose label was 00.
@@
The menu contained $101$ values of $\lambda$ between $0.2$ and $0.7$. We used the following success criteria: that the mean design label surpasses $\tau$, for $\tau\in[0,1.5]$ (for reference, the mean training label was $-4.8$), and that the exceedance over $1$ (i.e., the fraction of design labels that exceed $1$, using $g(Y)=\mathbf{1}[Y\geq 1]$) surpasses $\tau$, for $\tau\in[0,1]$ (for reference, the training labels’ exceedance over $1$ was $0.006$).
@@
We used the labels for all sequences in $\mathcal{X}$ (Wu et al.,, 2016) to compute the true value of $\theta_{\lambda}$ for all $\lambda\in\Lambda$. The error rate for each method was then computed as $\sum_{t=1}^{T}\mathbbm{1}[\exists\lambda\in\hat{\Lambda}\text{ s.t. }\theta_{\lambda}<\tau]/T$. We used $\alpha=0.1$ as a representative value for the desired error rate.
@@ Results
A good design algorithm selection method achieves a low error rate and a high selection rate for a variety of success criteria (settings of $g$ and $\tau$ in the criterion $\mathbb{E}_{Y\sim P_{Y;\lambda}}[g(Y)]\geq\tau$). For success criteria concerning the mean design label (i.e., $g$ is the identity), the prediction-only and CalibratedForecasts methods had error rates of 100% for most values of $\tau$ considered (Fig. 3a) Particularly for the prediction-only method, the mean design label produced by selected temperatures could be considerably lower than $\tau$ (Fig. 3b). In contrast, our method achieved error rates below the desired level of $\alpha$ for all values of $\tau$ considered. Its selection rate was 100% for a broad range of $\tau$, though it gradually declined for $\tau>1$, reflecting increasing conservativeness for these more stringent success criteria (Fig. 3a). We also ran our method with estimated density ratios for comparison, with similar results (see Appendix D.1 and Fig. 6 for details).
For GMMForecasts , recall that smaller values of the hyperparameter $q\in[0,1]$ yield forecasts that are, roughly speaking, more similar to the training data (see Appendix C.2 for details). Using $q\in\{0,0.5\}$ was prohibitively conservative and never selected anything on any trial. Using the maximum value, $q=1$, did yield high selection and low error rates for $\tau<0.5$. However, for greater values of $\tau$, the method ceased selecting anything, incorrectly indicating that no configuration could achieve these success criteria (Fig. 3a). All methods had similar qualitative performance for success criteria concerning the exceedance over $1$, though GMMForecasts was slightly less conservative (Fig. 7).
[…]
@@ 5.2 Algorithm Selection for Designing RNA Binders
These next experiments show the utility of our method in a setting that involves a variety of design algorithms and predictive models on the menu, and that requires density ratio estimation. The task is to design length-$50$ RNA sequences that bind well to an RNA target, where the label is the ViennaFold binding energy (Lorenz et al.,, 2011).
@@
To simulate training and held-out labeled data, we generated random mutants of a length-$50$ “seed” sequence, with $0.08$ probability of mutation at each site. Each mutant, $x$, was assigned a noisy label of its ViennaFold binding energy with the target: $y=\textsc{BindingEnergy}(x)+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma=0.02)$.
@@
The menu contained the following design algorithms and respective hyperparameter settings (see Appendix E.1 and Figs. 89 for details): AdaLead (Sinai et al.,, 2020), with its threshold hyperparameter, $\kappa$, set to different values in $[0.2,0.01]$; the algorithm used by Biswas et al., (2021), which approximately runs MCMC sampling from $p(x)\propto\exp(f(x)/T)$, with the temperature hyperparameter, $T$, set to values in $[0.005,0.02]$; Conditioning by Adaptive Sampling (Brookes et al.,, 2019) and Design by Adaptive Sampling (Brookes and Listgarten,, 2018), with their quantile hyperparameter, $Q$, set to values in $[0.1,0.9]$; and Proximal Exploration (Ren et al.,, 2022) with default hyperparameters, as the authors did not highlight any critical hyperparameters. Each of these design algorithms and respective hyperparameter settings was run with three different predictive models, resulting in a menu of $78$ configurations: a ridge regression model, an ensemble of fully connected neural networks, and an ensemble of convolutional neural networks. The success criterion was that the mean design label surpasses $\tau$, for $\tau\in[0.28,0.5]$ (for comparison, the mean training label was $0.28$). We also ran our method with an expanded menu of $249$ configurations involving additional hyperparameters, to assess how the multiple testing correction affects selection rates for larger menus (see Appendix E.1 and Fig. 8).
@@
Since the configurations on the menu did not have closed-form design densities, we used multinomial logistic regression-based density ratio estimation (MDRE) (Srivastava et al.,, 2023), which trains a classifier between designs and labeled sequences to estimate the density ratios between their distributions (see Appendix E.2 for details).